OpenStudy (anonymous):

Write an equation of a line parallel to y = 2/3x −5 in slope-intercept form that passes through the point (−6, −1).

OpenStudy (anonymous):

do you have a graph of that equation?

OpenStudy (anonymous):

Theres no graph but it has these answers y= 2/3x+3 y=2/3x+6 y= -3x-1 y=3x+6

OpenStudy (anonymous):

You know the slope is going to be 2/3 because it is parallel to the other line. Since you only have one point of the parallel line you need to first right it in point-slope form. Point-slope equation:y - y1 = m(x - x1) So for the parallel line this would y- (-1) = 2/3 (x- (-6)) can you transform this equation into slope- intercept?

OpenStudy (anonymous):

i know slope intercept form is y=mx+b but how would you change that to slope intercept?

OpenStudy (haseeb96):

hello first thing for the equation that is .It is parallel to the given equation so the slope of both equations should be same m1=m2 the slope of given equation is 2\3 m1=2\3 so m2 = 2\3 now use the point slope form and put the slope and given point in it (y- y1)=m(x -x1) y-(-1)=2\3(x-(-6)) y+1 = 2\3 (x + 6) 3(y+1) = 2(x+6) 3y+3 =2x+12 3y=2x+12 -3 3y=2x+9 for making slope intercept form put 3 on the other side y=2\3 x + 3 this is the correct answer

OpenStudy (anonymous):

First you want to simplify the equation so it would look like: y+1= 2/3 (x+6) distribute the 2/3: y+1= 2/3x + 4 you want to get y alone so subtract 1 from that side y = 2/3x + 3 so the answer is a!

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

No problem good luck with the rest of your assignment! :)